### v1

# CauSkelNet: Causal Representation Learning for Human Behaviour Analysis

Abstract
Constrained by the lack of model interpretability and a deep understanding of human movement in traditional movement recognition machine learning methods, this study introduces a novel representation learning method based on causal inference to better understand human joint dynamics and complex behaviors. We propose a two-stage framework that combines the Peter-Clark (PC) algorithm and Kullback-Leibler (KL) divergence to identify and quantify causal relationships between joints. Our method effectively captures interactions and produces interpretable, robust representations. Experiments on the EmoPain dataset show that our causal GCN outperforms traditional GCNs in accuracy, F1 score, and recall, especially in detecting protective behaviors. The model is also highly invariant to data scale changes, enhancing its reliability in practical applications. Our approach advances human motion analysis and paves the way for more adaptive intelligent healthcare solutions.
Index Terms:
Representation Learning, Human-Centered Computing, Causal Inference, Explainable AI, Personalized Healthcare
I Introduction
In recent decades, machine learning has made significant strides in the field of sports science, particularly in behavioural analysis and performance evaluation [1][2][3]. The integrated application of computer-aided data processing, mathematical modelling, and simulation optimisation has accelerated the development of behavioural analysis methods and broadened the scope of sports technique exploration [4][5]. These advancements have enabled in-depth analysis of complex behavioural datasets, which was previously challenging using traditional methods [6].
Human behaviour is a multidimensional, complex phenomenon encompassing physiological and psychological aspects, profoundly influenced by emotions and cognition [7]. Research indicates that interactions between behaviours involve intricate Bayesian causal frameworks, which are further modulated by emotional states [7][8]. This complexity necessitates the adoption of more advanced, interpretable algorithmic approaches to comprehensively understand and simulate these relationships.
Despite substantial progress in movement recognition and prediction through machine learning, contemporary research often overlooks rigorous exploration of the underlying causal relationships behind behavioural dynamics. This limitation leads to insufficient model interpretability, thereby impacting our in-depth understanding of human behaviour. Such oversight not only restricts the explanatory power of movement recognition models but also diminishes their potential in guiding optimisation and intervention strategies [6]. For instance, in the field of gait analysis, although machine learning has been widely applied, most studies have yet to delve into the fundamental causal relationships embedded within biomechanical data [9].
To overcome the limitations of existing research, this work proposes an innovative approach combining KL divergence with causal inference, aimed at simulating the complex interactions between human joints in behavioural contexts, thereby learning more effective representations of human kinematic data. We integrate this method with graph convolutional neural networks, determining the efficacy of learned representations through comparison with baseline models. This approach not only pioneers a new application of causal inference learning in the field of human kinematics but also provides a mathematically rigorous framework for elucidating the dynamic interrelationships between joints. This novel representation learning method can reveal joint interactions across different behavioural patterns, deepening our understanding of overall movement patterns and paving the way for new research directions in human-centred computation and behavioural analysis. The core contributions of this research are as follows:
  • Innovative Causal Representation Learning Framework: The study introduces a novel framework that integrates causal inference into human motion analysis. This framework successfully captures the causal relationships between joints, providing a new perspective on representing human movement.
  • Combination of PC Algorithm and KL Divergence: The research combines the PC algorithm with KL divergence in a two-stage causal representation learning method. The PC algorithm is used to uncover potential causal structures, while KL divergence quantifies the strength of mutual causal relationships, resulting in more precise and meaningful motion representations.
  • Highly Interpretable and Robust Motion Recognition: The study develops a motion recognition representation learning method that is both highly interpretable and robust. This approach generates easily comprehensible causal relationship graphs and demonstrates strong invariance to data scale changes, enhancing the model’s reliability and applicability in practical scenarios.
II Related Work
II-A Traditional Kinesiology Research
Traditional kinematic studies primarily employ motion capture and bio-mechanical modelling to analyse human movement mechanisms. These methods quantify the dynamic changes of various body parts during movement, revealing the complex interactions between physical activities and physiological functions [10][11]. However, these approaches have inherent limitations in representation learning. They mainly quantify observable physical indicators (such as joint angles and force output) but fail to effectively learn and represent complex inter-joint relationships. For instance, James et al. [10] demonstrated that fatigue significantly altered bio-mechanical and neuromuscular responses during landing, highlighting the inadequacies of traditional bio-mechanical frameworks in capturing and representing factors of performance variation. Furthermore, Cop et al.[12] found that while the anatomical and functional characteristics of musculoskeletal structures are well understood, the causal interactions between neural and bio-mechanical levels in specific movements remain difficult to represent effectively. These complex physiological and neuromechanical dynamics suggest that traditional methods struggle to comprehensively learn and represent the multidimensional factors influencing motor performance. This underscores the necessity of introducing more advanced representation learning methods in kinematic studies to better capture and understand the intrinsic mechanisms of human movement.
II-B Multi-modal Data and Representation Learning
Advancements in deep learning and multi-modal data modelling have significantly enhanced our understanding of human behaviour. Researchers have developed more sophisticated and effective behavioural representation methods by utilising various data sources, including motion capture, bio-mechanical analysis, and physiological monitoring. The implementation of Graph Convolution Networks (GCN) and Parallel Spatiotemporal Encoding Motion Representation (P-STEMR) models has paved new avenues for analysing movement dynamics [2]. GCN optimises joint data processing, effectively learning and representing inter-joint interactions, and exploring potential pathways influencing movement dynamics. Concurrently, the Human Activity Recognition-Protective Behaviour Detection (HAR-PBD) framework has enhanced the ability to represent complex behavioural patterns through continuous monitoring of joint dynamics and interactions [13]. However, existing research primarily focuses on unidirectional influences, overlooking the complex interplay and causal relationships between behaviour, emotions, and other factors. This limitation underscores the need for more in-depth causal analysis to develop advanced learning models capable of comprehensively representing the complexity of human behaviour. Such an approach can not only improve the accuracy of behaviour recognition but also provide a richer informational basis for developing personalised intervention strategies.
III Methodology
To address the research gap in causal relationships among joints during specific behaviors, we employ causal graphical models [14, 15, 16, 17] to explore causal structures within the human anatomical framework. We use Directed Acyclic Graphs (DAGs) with nodes representing joints and edges denoting causal interactions [14]. The graph’s topology is based on human physiology, with nodes at critical joints and edges mapping physical links. This approach captures the inherent connectivity among body segments and provides a foundation for detailed causal investigations into human pain responses and behavior.
Specifically, we employ the PC algorithm and Kullback-Leibler (KL) divergence to develop and validate causal graphical models, analyzing interactions among human joints[18, 19]. The PC algorithm identifies causal structures through conditional independence tests as:
$X\perp\!\!\!\perp Y\mid S$ (1)
This algorithm is enhanced to establish a preliminary undirected graph, capturing latent connections among variables[20, 21, 22]. To determine causality directions, we employ Kullback-Leibler (KL) divergence, which quantifies the difference between probability distributions P and Q based on following Equation3:
$D_{\text{KL}}(P\parallel Q)=\sum_{x\in\mathcal{X}}P(x)\log\left(\frac{P(x)}{Q(x)}\right)$ (2)
To assess the influence strength between nodes and infer causal directions, for any two nodes $(i)$ and $(j)$, we calculate the KL divergence between $P(j\mid i)$ and $P(j)$, and vice versa. This reveals asymmetry in information flow, indicating potential causal directions. A larger KL divergence suggests reduced prediction uncertainty of node $(j)$ given node $(i)$, implying a likely causal direction. This method identifies direct connections and quantifies causal strengths, providing a robust foundation for establishing causal directions in DAGs. It extends to dynamically changing biomedical data, elucidating intricate interactions between movement behaviors and pain responses.
However, KL divergence has limitations: Firstly the sensitivity to sample size potentially affects causal direction determination. Next, as a non-symmetric metric, it illustrates improved prediction accuracy rather than direct causal strength. To address these issues, we employ statistical corrections and comparative analysis with additional techniques, namely information criteria and cross-validation, to refine causal direction assessments. These measures ensure more reliable causal inferences across different models and datasets.
III-A Formulation
Let $\mathcal{G}=(V,E)$ be a graph representing the human body structure, where $V=\{v_{1},\ldots,v_{n}\}$ denotes joints and $E$ denotes connections between joints. Our objective is to derive a representation $\phi:V\rightarrow\mathbb{R}^{d}$ that encapsulates the causal relationships between joints.
We employ the PC algorithm to elucidate the causal structure. For variables $X,Y\in V$ and a set of conditioning variables $S\subset V\setminus\{X,Y\}$, we examine the conditional independence with Eq.1. The PC algorithm yields a completed partially directed acyclic graph (CPDAG) $\mathcal{G}^{*}=(V,E^{*})$.
To quantify the strength of causal relationships and derive joint representations, we utilize KL divergence. For any edge $(i,j)\in E^{*}$, we define:
$D_{KL}(i\rightarrow j)=\sum_{x_{j}\in\mathcal{X}_{j}}P(x_{j}\mid x_{i})\log\left(\frac{P(x_{j}\mid x_{i})}{P(x_{j})}\right)$ (3)
Subsequently, we formulate a representation function $\phi$ as:
$\phi(v_{i})=\left[\frac{D_{KL}(i\rightarrow j)}{\sum_{k\in\text{Ne}(i)}D_{KL}(i\rightarrow k)}\right]_{j\in\text{Ne}(i)}$ (4)
where $\text{Ne}(i)$ denotes the set of neighbors of node $i$ in $\mathcal{G}^{*}$.
III-B Theoretical Guarantees
Theorem 1.
The representation $\phi$ preserves the causal structure of $\mathcal{G}^{*}$. (See Proof in Appendix A-A)
Theorem 2.
The representation $\phi$ is invariant to the scale of the data. (See Proof in Appendix A-B)
Algorithm 1 Causal Analysis Using PC Algorithm and KL Divergence
1:Initialization: Construct graph skeleton $G=(V,E)$ from data $\mathcal{D}$ using PC algorithm.
2:Refinement: Refine $G$ by adjusting edges $\epsilon\in E$ based on additional conditional independence tests.
3:Bayesian Network Formation: Formulate DAG $\mathcal{G}=(V,\vec{E})$ by directing $E$ using dependencies derived from $\mathcal{D}$.
4:KL Divergence Calculation:
5:      Estimate conditional probabilities $P(j|i)$ for all $i,j\in V$ using Maximum Likelihood Estimation.
6:      Estimate marginal probabilities $P(j)$ for all $j\in V$.
7:      For each pair $(i,j)\in\vec{E}$, compute:
8:           $D_{\text{KL}}(i\rightarrow j)=\sum_{x_{j}\in\mathcal{X}_{j}}P(x_{j}|x_{i})\log\left(\frac{P(x_{j}|x_{i})}{P(x_{j})}\right)$
9:           $D_{\text{KL}}(j\rightarrow i)=\sum_{x_{i}\in\mathcal{X}_{i}}P(x_{i}|x_{j})\log\left(\frac{P(x_{i}|x_{j})}{P(x_{i})}\right)$
10:Causal Direction Assessment: Determine the direction of causality based on the asymmetry of the KL divergence values. A higher $D_{\text{KL}}$ from $i$ to $j$ compared to from $j$ to $i$ suggests that $i$ causally influences $j$.
11:Edge Weight Assignment: Assign weights to the edges in the DAG based on the computed KL divergence values, reflecting the strength of the causal influence.
12:Model Validation: Use statistical tests and cross-validation techniques to validate the causal model, ensuring robustness and reliability of the inferred causal relationships.
IV Experiment
IV-A Dataset
The EmoPain dataset [23] employed in this study is designed to analyze movement behaviors and pain recognition in chronic pain (CP) patients relative to a healthy control group [24]. This dataset records pain levels and protective behaviors from 18 IMUs of 12 healthy and 18 CP participants within a controlled movement challenge, providing insights into how pain influences movement. As depicted in Figure 1, the dataset constructs a fundamental human causal graph using the X, Y, and Z coordinates of 22 body joints, capturing body postures during movements and facilitating in-depth analyses of human motion patterns.
Figure 1: Joint number and relative information in EmoPain Dataset
IV-B Experiment Design
In this study, depicted in Figure 1, we used the EmoPain dataset to create an undirected causal graph model of human structure, examining causal connections between joints in protective and non-protective behaviors and their impact on pain analysis. We processed the three-dimensional joint coordinate data (X, Y, Z) to capture dynamic interactions, then applied the PC algorithm to learn a DAG of potential causal relationships. A Bayesian network was constructed, and conditional probability distributions were determined using maximum likelihood estimation. For each joint pair (e.g., Joint 1 and Joint 2), we calculated the conditional probability distribution and used KL divergence to measure information transfer efficiency, reflecting the reduction in uncertainty of one joint’s state given another’s. We compared KL divergence across dimensions to identify the dominant causal direction, selecting the most frequent direction and using its average KL divergence as a weight for information transfer efficiency. This analysis delved into joint interactions and their role in pain perception and protective behaviors, offering new insights for behavior-based pain analysis.
IV-C Experiment Analysis
Figure 2: The two different types of casual graphs for protective and non-protective behaviors. Some differences between the models could be found: e.g. edge between joint 9 and 20 and the edge weight
Figure 3: Causal graphs for four different exercise types in both protective and non-protective behaviors are displayed, from left to right: One leg stand, Reach forward, Bend, and Others. Each colored box highlights the differences between protective and non-protective behaviors based on graph representations. For instance, the direction of the arrows are different.
Causal graphs for protective and non-protective behaviors were plotted using 30,000 data samples, with Kullback-Leibler divergence normalized for comparison. Dashed lines indicate imprecise causal connections. As shown in Figure 2, significant differences between protective and non-protective behaviors were observed, such as reversed causal relationships and varied weights between certain nodes (e.g., nodes 9 and 20). These differences suggest that the human body adopts various physical adjustment strategies in response to pain. For instance, protective behaviors may restrict joint movements, while non-protective behaviors allow wider motion ranges. Similar variations were observed between other node pairs under different pain perceptions and emotional states. These findings provide insights into how pain influence human motor behavior, offering a bio-mechanical basis for personalized medical and adaptive machine learning models. In practical applications, this allows for more precise therapeutic recommendations and movement guidance based on individual behavior types.
To explore variations in joint interactions between protective and non-protective behaviors under different movement types, we plotted causal graph models labeled by exercise type, as shown in Figure 3. By comparing causal graphs across different behavioral contexts, we observed consistency in the causal relationships among joints under similar pain perceptions and protective mechanisms. However, significant differences emerged when pain triggered a shift from non-protective to protective behaviors, demonstrating how pain perception and protective behaviors jointly influence human motion. These findings validate our hypothesis: under pain and emotional distress stimuli, causal connections among joints undergo notable changes as protective behaviors are prompted. In protective behaviors, to alleviate pain or prevent further injury, the activity of certain joints can be consciously restricted, reflected by changes in the direction and strength of causal relationships. Conversely, in non-protective behaviors, joints may exhibit greater freedom of movement, displaying different neural responses and behavioral adaptation to pain, with weaker causal constraints.
TABLE I: Performance Metrics for Baseline vs. Causal GCN in test
Exercise Type Baseline Causal GCN
Acc. Prec. F1 Rec. Acc. Prec. F1 Rec.
One Leg Stand 0.933 0.895 0.862 0.910 0.942 0.903 0.886 0.969
Bend 0.883 0.894 0.859 0.792 0.916 0.915 0.908 0.920
Reach Forward 0.882 0.899 0.826 0.736 0.892 0.897 0.849 0.794
Other 0.939 0.924 0.474 0.320 0.940 0.878 0.569 0.439
Figure 4: Comparative Trend Analysis of Performance Metrics Across Four Exercise Types for Baseline and Causal GCN Models.
V Evaluation
V-A Evaluation Design
In this study, we built upon our previous research involving causal directed acyclic graphs (DAGs). Our preliminary results showed that the causal graph model effectively captured changes in joint connections under different behavioral states. To advance this, we designed a comprehensive evaluation approach:
  • Baseline Model: We used an undirected graph based on human anatomy as our baseline, inspired by the success of Graph Convolutional Networks (GCN) in similar datasets [13]. (See GCN model detail in Fig 5)
  • Comparative Analysis: We compared causal graph models of protective behaviors across various movement types with the baseline model, maintaining interpretability by avoiding edge weight additions.
  • Generalizability Test: We trained and evaluated a generalized causal model using 30,000 data samples to test its performance across different movement types.
This evaluation strategy aims to assess the practical performance and adaptability of causal graph models in complex behavior recognition tasks. It not only advances causal analysis methods for behavior and emotion computation but also provides interpretable foundations for understanding human movement.
Figure 5: Architecture of the Graph Convolutional Network (GCN) Model
V-B Evaluation and Result
As shown in Table I, Causal GCN outperformed baseline models across various movement types. For single-leg standing, Causal GCN achieved superior F1 score (0.886) and recall (0.969) compared to the baseline (0.862 and 0.910 respectively). In ”bending” activities, it improved recall from 0.792 to 0.920. For complex ”other” movements, F1 score increased from 0.474 to 0.569 and recall from 0.320 to 0.439.
Figure 4 illustrates Causal GCN’s consistent performance advantage across all movement types. This superiority in accuracy, F1 score, and recall demonstrates its effectiveness in capturing subtle joint dynamics and behavioral relationships.
These results highlight the benefits of integrating causal structures into behavioral analysis models, enhancing both interpretability and performance. The Causal GCN’s adaptability to different movement types and its ability to capture behavior changes associated with pain make it promising for developing personalized treatment strategies.
Figure 6: Comparative Evaluation of Baseline and Causal GCN Performance Metrics Without Exercise Type Labeling
TABLE II: Performance Metrics Comparison of General Models without Exercise Type Labels
Model Accuracy Precision F1 Recall
Baseline - Test 0.928 0.900 0.724 0.625
Causal GCN - Test 0.942 0.908 0.793 0.733
To assess the generalization capability of causal graph models in behavior recognition, we compared the performance of the baseline model and the general causal graph model without exercise type labels. As shown in Table II, the Causal GCN model outperformed the baseline across all metrics over 50 epochs of evaluation. Causal GCN significantly improved performance metrics compared to the baseline model: testing F1 score (0.793 vs. 0.724), and testing recall (0.733 vs. 0.625). Figure 6 illustrates the Causal GCN’s superior and more stable performance across all indicators. These results demonstrate the model’s enhanced capability in identifying protective behaviors and its adaptability across different behavior types, confirming that integrating causal structures improves both generalizability and interpretability in behavior recognition tasks.
VI Discussion
This study introduces an innovative causal representation learning approach, advancing the field of human motion and behavior analysis. By synergizing causal inference with state-of-the-art representation learning techniques, we propose a novel dimensional perspective for information extraction, thereby enhancing model performance, interpretability, and generalization capabilities. This methodology demonstrates significant efficacy in feature extraction from complex human data, substantially improving the performance and interpretability of human-centered computational models. Our approach underscores the critical role of causal reasoning in representation learning, emphasizing that elucidating deeper feature relationships beyond mere statistical correlations is paramount for comprehending intricate human behavior patterns[25]. The findings suggest that capturing causal structures in data can lead to more robust and generalizable models, particularly in the context of human behavior analysis. Future research directions may include the integration of more sophisticated causal discovery algorithms into the representation learning framework, potentially revolutionizing the analysis of complex human data across various domains, including healthcare, human-computer interaction, and affective computing.
Furthermore, the proposed causal representation learning method shows potential advantages in analyzing the differential causal relationships between protective and non-protective joint behaviors, with findings aligning with existing medical literature on the impact of pain and emotional distress on behavior. Through the analysis of the EmoPain dataset, this study identified behavior patterns consistent with several established medical models: in line with Vlaeyen and Linton’s fear-avoidance model, chronic low back pain patients exhibited stronger causal relationships between hip and knee joints during bending tasks[26]; similarly, consistent with Sullivan et al.’s pain catastrophizing theory, participants with high pain catastrophizing showed increased causal strength between trunk and lower limb joints during sit-to-stand tasks[27]. Additionally, supporting Hodges and Tucker’s motor adaptation model, chronic pain patients demonstrated more complex causal networks among spinal joints during extension movements[28]. This method also provides a novel perspective for identifying pain-related fear-avoidance behaviors as described by Asmundson et al. , and by integrating emotional score data, it captures the interaction between emotional states and chronic pain emphasized by Geisser et al., such as the generalized reduction in causal joint strength observed in participants with high depression levels[29][30]. These findings not only offer a new data-driven approach for validating existing medical theories but also provide potential applications for quantifying emotion-behavior relationships and developing personalized pain management strategies.
VII Conclusion
In this study, we propose an innovative causal representation learning method aimed at deepening our understanding of human movement and behavioural patterns. Utilising the EmoPain dataset, we have developed a two-stage framework combining the Peter-Clark algorithm and KL divergence to learn representations of complex causal relationships between human joints[31][18].
This model integrates biomechanical representation extraction into human motion analysis, enhancing machine learning models’ interpretability and performance. By capturing inter-joint causal relationships, our method accurately identifies and responds to pain-induced motion behavior changes. The causal GCN model outperforms the baseline undirected GCN across multiple metrics, demonstrating its efficacy and applicability. Our approach generates highly interpretable and robust representations, surpassing traditional graph convolutional neural networks in identifying pain-induced behavioral changes. It excels in handling emotion-influenced protective behaviors and shows strong generalization across different movement types, enhancing practical robustness. The innovation lies in capturing subtle causal relationships overlooked by traditional methods, providing deeper insights into human behavior analysis. This causal representation learning method offers a novel approach to understanding complex human motion patterns and their underlying mechanisms.
Our research not only advances the integration of machine learning and personalised medicine but also opens new avenues for personalised adaptive healthcare and behavioural analysis. By introducing causal inference into representation learning, our method lays the foundation for developing more precise and adaptive intelligent healthcare solutions. This approach holds immense potential in improving the interpretability and personalised adaptability of human-centred computational models.
Looking ahead, this research prides new ideas for constructing more intelligent and humanised medical systems. It not only helps improve the accuracy of personalised behavioural analysis but may also find broad applications in related fields such as affective computing and rehabilitation medicine. We believe that this representation learning method, integrating causal inference, will open new prospects for the application of artificial intelligence in healthcare, driving further development of personalised medicine.
Appendix A Detailed Proof of Theorem
A-A Proof of Theorem 1
Proof.
Let $(i,j)\in E^{*}$ be an edge in the CPDAG $\mathcal{G}^{*}$ output by the PC algorithm.
If there is a causal relationship from $i$ to $j$:
  • a)
    The PC algorithm will not remove this edge during the conditional independence tests.
  • b)
    The KL divergence $D_{KL}(i\rightarrow j)$ will be positive:
    $D_{KL}(i\rightarrow j)=\sum_{x_{j}\in\mathcal{X}_{j}}P(x_{j}\mid x_{i})\log\left(\frac{P(x_{j}\mid x_{i})}{P(x_{j})}\right)>0$
  • c)
    Therefore, in the representation $\phi(v_{i})$:
    $\phi(v_{i})_{j}=\frac{D_{KL}(i\rightarrow j)}{\sum_{k\in\text{Ne}(i)}D_{KL}(i\rightarrow k)}>0$
Conversely, if there is no causal relationship from $i$ to $j$:
  • a)
    The PC algorithm will remove this edge during the conditional independence tests.
  • b)
    Since $j\notin\text{Ne}(i)$, $\phi(v_{i})_{j}$ will be undefined (effectively zero).
Thus:
$\phi(v_{i})_{j}>0\iff(i,j)\in\mathcal{G}^{*}$
A-B Proof of Theorem 2
Proof.
Let $\alpha>0$ be a scaling factor applied to the data.
For any two probability distributions $P$ and $Q$, KL divergence is invariant under scale transformations:
$D_{KL}(\alpha P||\alpha Q)=\sum_{x}\alpha P(x)\log\left(\frac{\alpha P(x)}{\alpha Q(x)}\right)$
$=\alpha\sum_{x}P(x)\log\left(\frac{P(x)}{Q(x)}\right)$
$=\alpha D_{KL}(P||Q)$
In our representation $\phi$:
$\phi(v_{i})_{j}=\frac{D_{KL}(i\rightarrow j)}{\sum_{k\in\text{Ne}(i)}D_{KL}(i\rightarrow k)}$
Under a scale transformation by $\alpha$:
$\phi^{\prime}(v_{i})_{j}=\frac{\alpha D_{KL}(i\rightarrow j)}{\sum_{k\in\text{Ne}(i)}\alpha D_{KL}(i\rightarrow k)}$
$=\frac{\alpha D_{KL}(i\rightarrow j)}{\alpha\sum_{k\in\text{Ne}(i)}D_{KL}(i\rightarrow k)}$
$=\frac{D_{KL}(i\rightarrow j)}{\sum_{k\in\text{Ne}(i)}D_{KL}(i\rightarrow k)}$
$=\phi(v_{i})_{j}$
Therefore, $\phi^{\prime}(v_{i})_{j}=\phi(v_{i})_{j}$ for all $i,j$, proving that $\phi$ is invariant to scale transformations of the data.